A firm hires labor and capital to produce tomatoes. Currently the marginal product of the last unit of labor input is 20 and the marginal product of the last unit of capital input is 30. The market wage for labor is $10 a. If the firm is using the optimal combination of inputs, then the price of capital is?

Select one:

a. $15

b. $50

c. $3

d. $20
Answer #1

ANswer

Optimum input combination is where the per dollar marginal product from each input is equal and that is depicted as

MPL/MPK=w/r

or

MPL/w=MPK/r

w=wage, r=price of capital

r=(MPK*w)/MPL

r=(30*10)/20

r=15

the price of capital is r=$15

Option a